Two spheres are launched horizontally from a 1.14-m-high table. Sphere A is launched with an initial speed of 5.77 m/s. Sphere B is launched with an initial speed of 2.34 m/s.

A) What is the time for first sphere to hit the floor?

B) What is the time for second sphere to hit the floor?


C) What is the distance that the first sphere travels from the edge of the table?

D)What is the distance that the second sphere travels from the edge of the table?
Answer #1

a)

time taken by 1st sphere  

d = 0.5 g t1^2

1.14 = 4.9* t1^2

t1 = 0.482 s

time taken by 2nd sphere

1.14 = 0.5* 9.8* t2^2

t2 = 0.482 s

so both the spheres reach ground simultaneously.

=====

b)

t2 = 0.482 second

=====

c)

d1 = speed1* t1 = 5.77* 0.482 = 2.781 m

======

d)

d2 = speed2 * t2 = 2.34* 0.482 = 1.129 m

======

do comment in case any doubt, will reply for sure.. Goodluck